You have a 240-Ω resistor and a 0.360-H inductor. Suppose you take the resistor and inductor and make a series circuit with a voltage source that has a voltage amplitude of 32.0 V and an angular frequency of 270 rad/s .

What is the impedance of the circuit?

What is the current amplitude?

What is the voltage amplitude across the resistor?

What is the voltage amplitudes across the inductor?

What is the phase angle ϕ of the source voltage with respect to the current?
Answer #1

Inductive reactance, XL = ωL = 270 x 0.36 = 97.2 ohm

A)

Impedance, Z = sqrt(240^2 + 97.2^2) = 258.94 ohm

B)

Current, I = V/Z = 32/258.94 = 0.124 A

C)

Voltage, Vr = IR = 0.124 x 240 = 29.66 V

D)

Voltage, V = I Xl = 0.124 x 97.2 = 12.01 V

E)

Phase angle, phi = arctan(97.2/240) = 22.05°

Comment in case any doubt please rate my answer....